OpenStudy (adunb8):

this is differential equation & linear algebra question: im trying to looking for exact equation (x^2-y)y'+2x^3+2xy=0 the answer needs to be 3.) y=x^2 +_ (2x^4+c)^1/2

OpenStudy (amistre64):

integrate one side up, add a generic constant function to it; and derivative it down wrt the other variable. Then compare to find the derivative of the generic function

OpenStudy (adunb8):

can you steps? i cant really see it with just writing

OpenStudy (amistre64):

\[(x^2-y)dy+(2x^3+2xy)dx=0\] \[\int (x^2-y)dy=x^2y-\frac12y^2+g(x)\] \[\frac{d}{dx}(x^2y-\frac12y^2+g(x))=2xy+g'(x)\] \[\cancel{2xy}+g'(x)=2x^3+\cancel{2xy}\]

OpenStudy (amistre64):

integrate g' to determine g and fill it in; then check the work :)

OpenStudy (adunb8):

here what i got i did My = 2x^3 + 2xy then did partial derivate respect to y then got 2x then Nx = (x^2-y) and got 2x also which is same so we can conclude that it is exact equation and then i did Mdx = 2x^3 +2xy which gives 1/2x^4+x^2y + g'(y)

OpenStudy (amistre64):

why did you do M wrt y?

OpenStudy (adunb8):

M(x,y) = N(x,y) =

OpenStudy (amistre64):

if we apply your workings \[(x^2-y)y'+2x^3+2xy=0\] \[(x^2-y)\frac{dy}{dx}+2x^3+2xy=0\] \[(x^2-y)dy+(2x^3+2xy)dx=0\] \[Ndy+Mdx=0\] M is not integrated wrt y

OpenStudy (amistre64):

i read you wrong :) you were proving it was exact

OpenStudy (adunb8):

yeah

OpenStudy (adunb8):

and then i did integrate to Mx

OpenStudy (amistre64):

it doesnt matter which side you start integrating; i started on the "N" in this case

OpenStudy (adunb8):

okay but i dont seem to get the answer provided by my teacher

OpenStudy (amistre64):

your setup is going along fine; except, its not g'(y); its just a generic function of y: g(y)

OpenStudy (adunb8):

oh okay so then after i did Mx i now integrate it to respect to Ndy then im stuck...

OpenStudy (amistre64):

it tends to be easier if you take the derivative down to Ndy to compare it with

OpenStudy (adunb8):

instead of using Mdx first?

OpenStudy (amistre64):

you already did the "first"

OpenStudy (amistre64):